OpenStudy (anonymous):

Let f(x) = x^2 – 81. Find f–1(x). Can someone double-check what I have?

OpenStudy (hba):

Show me what ya got :)

OpenStudy (anonymous):

Okay so I know first to substitute f(x) with y and then to reverse the x & y variables which would give me x=y^2-81. Then I use a root on both sides and get \[\pm x=y-9\] and finally I add 9 on both sides to get \[\pm 9 \sqrt{x} = y\]

hartnn (hartnn):

how u got y- 9 ??

OpenStudy (hba):

Interesting :)

hartnn (hartnn):

but incorrect :P

OpenStudy (anonymous):

well we have y^2-81, idk factoring it would be (y+9)(y-9) but am I supposed to do that?

hartnn (hartnn):

no, what you did was \(\sqrt{y^2-81}=y-9\) that is incorrect. whats to be done : let y= x^2-81 add 81 to both sides. then take square root of both sides...

OpenStudy (anonymous):

oh okay :) so it'd be x+81= y^2 and once we square root both sides would the answer come out to \[\pm 9 \sqrt{x}=y\]

OpenStudy (hba):

@hartnn Bache mai halka haat rakho :)

hartnn (hartnn):

taking square root on both sides of \(x+81=y^2\) will give you \(\sqrt{x+81}=\sqrt{y^2}=y\) to get the inverse function as \(\sqrt{x+81}\) got this ? hba, i didn't get you.

OpenStudy (anonymous):

okay I get it :) thx a ton! :)

hartnn (hartnn):

welcome ^_^